the width of a rectangle is 8 inches more than the length. The perimeter is 32. find the length and width of the rectangle.
Answer by Cromlix(4381)

Length = x
Width = x + 8
Perimeter = 2* Length + 2* Width
32 = 2*x + 2*(x + 8)
32 = 2x + 2x + 16
32 = 4x + 16
32 - 16 = 4x
4x = 16
x = 4
Length = 4ins.
Width = 12ins.
Hope this helps.
:-)